A man wants to build a teepee for his kids. He plans to use 10 foot poles for the supports and wants the base of the teepee to be 12 foot across. What would the height be?
It has been years since I did anything like this so I don't even know where to begin. It's an algebra problem.

Answer by stanbon(75887)

A man wants to build a teepee for his kids. He plans to use 10 foot poles for the supports and wants the base of the teepee to be 12 foot across. What would the height be?
----
The height is the pole in the middle of the teepee.
---
Let the height be "h".
---
From the bottom of the height-pole to one of the support poles
is (1/2) of the base = 6 ft.
----
The support pole extends on a slant from the top of the height-pole
to the outer end of the base.
-----
If you draw those three pieces you will have a right triangle
with base = 6 , height = "h" , and hypotenuse = 10ft.
-----
Using Pythagoras you can solve for "h":
10^2 = h^2 + 6^2
h^2 = 100 - 36
h^2 = 64
height = 8 ft.
=====================
Cheers,
Stan H.